# Automatically Analyzing Performance Issues in Android Apps: How Far Are We?

### v2

@@ 3.2.2. Result - Technique
Lyu et al. (Lyu2018RemoveRF, ) 2018 206 Real-World Energy Consumption Responsiveness Database Operations Loop  (Lyu2018RemoveRFTool, )
Vásquez et al. (LinaresVsquez2018MultiObjectiveOO, ) 2018 5 Real-World Energy Consumption Display Screen  (LinaresVsquez2018MultiObjectiveOOTool, )
Habchi et al. (Habchi2019TheRO, ) 2019 324 Real-World Memoy Consumption Energy Consumption Responsiveness UI Rendering View Inefficient Method  (Habchi2019TheROData, )
Li et al. (Li2019CharacterizingAD, ) 2019 19 Real-World Memory Consumption Energy Consumption Responsiveness Image  (Li2019CharacterizingADTool, )
AsyncBench (Pan2020StaticAC, ) 2020 69 Micro Memory Consumption Energy Consumption Responsiveness AsyncTask  (Pan2020StaticACTool, )
Liu et al. (Liu2019DroidLeaksAC, ) 2020 34 Real-World Memory Consumption Energy Consumption Resource  (Liu2019DroidLeaksACTool, )
Das et al. (Das2020CharacterizingTE, ) 2020 316 Real-World Memory Consumption Energy Consumption UI Rendering Data Type Inefficient Method View Wakelock  (Das2020CharacterizingTETool, )
Lu et al. (Lu2022DetectingRU, ) 2022 35 Real-World Memory Consumption Energy Consumption Resource  (Lu2022DetectingRUTool, )
Janssen et al. (Janssen2022OnTI, ) 2022 40 Real-World Energy Consumption Responsiveness UI Rendering  (Janssen2022OnTITool, )
Petalotis et al. (Petalotis2024AnES, ) 2024 9 Real-World Energy Consumption Responsiveness Advertisements  (Petalotis2024AnESTool, )
@@ 3.2.3. Result - Dataset
To collect datasets, several types of papers need to be included: 1) Tool papers that construct datasets to evaluate their designed tools; 2) Empirical studies that collect specific datasets to analyze performance issues, where labeled data can be used to support evaluation. After examining all the public links or websites within the 66 identified studies, we identified 10 (15.15%) unique, available, and open-sourced datasets, as shown in Table 4. Notably, we conducted the examination in October 2024. Since most of these datasets are designed for diverse performance factors, it becomes infeasible to ensure a fair comparison. In addition, the number of factors covered by these datasets is quite limited. Out of the 72 factors identified in our taxonomy, only 24 (33.33%) are addressed by existing datasets. This indicates a significant need for constructing more comprehensive datasets. Moreover, there is a considerable imbalance in dataset size. The smallest dataset contains only five apps, while the largest has 324, with an average of 105.7 apps per dataset. This imbalance makes it unrealistic for any single dataset to cover all potential issues comprehensively. Even if some tools are effective at detecting performance issues, they may still face soundness challenges due to false negatives. Therefore, more effort is needed to create datasets that are not only comprehensive but also more representative of the diverse range of performance issues.
Answers to RQ2.3: We identified 10 open-source datasets for evaluating performance issues; however, they encompass only 24 out of 72 factors (33.33%), highlighting a significant gap in tool evaluation. These datasets suffer from inadequate sample sizes, with the smallest containing only five apps and the largest consisting of 324.
@@ 4.1. Targeted Objective Imbalance
In Section 3.2.1, we found that there is a misalignment in the primary concerns of researchers, developers, and users: researchers primarily focus on energy consumption, developers prioritize memory consumption, while users are most concerned with responsiveness. This suggests that future research should increase focus on responsiveness and memory issues, as these have a more immediate impact on both end-users and developers. We also found two categories of performance issues—storage consumption and internet data usage—that remain unexplored in academic research, indicating potential areas for further investigation. Additionally, only 42.86% of the factors contributing to performance issues observed in real-world settings have been explored in academic research, remaining nearly 60% still awaiting for researchers to investigate. Interestingly, there are 23 factors among the 72 factors lead to performance issues are only been explored in literature but not observed in the real-world scenarios. This suggests that the issues currently discussed in academic research may not be of real concern to users or developers in real-world settings. Overall, this imbalance between academic research objectives and real-world challenges underscores the need for greater focus on the real-world performance issues and contributing factors that are most relevant to end-users and developers.
@@ 4.2. Automatic Performance Analysis Tools are Still in an Early Stage
Our study revealed that only 17 out of 72 factors (23.61%) are addressed by state-of-the-art (SOTA) tools, with most tools focusing on detection rather than resolution. In fact, only two out of the existing tools are dedicated to fixing performance issues. Although static, dynamic, and deep learning (DL) techniques have shown effectiveness in research experiments, applying these methods to holistically detect and fix performance issues in practice remains an unsolved challenge. Future research should focus not only on building comprehensive frameworks to detect real-world performance issues, but also on providing solutions for fixing these issues.
The lack of variety in technique selection hinders the soundness, as static approaches are not proficient in handling complicated code features (e.g., reflection (sun2021taming, ), obfuscation (samhi2022jucify, ), and hardening (Schloegel2022LokiHC, )), and it’s non-trivial for dynamic tools to achieve satisfactory results to trigger sufficient paths in practice. Thus, developing effective tools against performance issues will be a noteworthy research topic in the future. In fact, DL has made considerable achievements in software engineering (Wang2023MachineDeepLF, ; Feng2020CodeBERTAP, ; Gu2016DeepAL, ; Hu2018DeepCC, ), and these advanced techniques have demonstrated powerful capabilities in solving many complex tasks (i.e. code analysis (White2016DeepLC, ; Jebnoun2021ClonesID, ; Zhao2018DeepSimDL, ) and code generation (Wang2022TestDrivenML, )). Applying advanced DL techniques to detect and fix Android performance issues is promising, although significant challenges remain in collecting sufficient performance issue dataset for model training. The emergence of large language models (LLMs) also presents new possibilities. LLMs, with their in-context learning capabilities, could address the limitations of traditional DL approaches by enabling models to detect and resolve performance issues with fewer examples and greater adaptability (Hou2023LargeLM, ; Liao2023A3CodGenAR, ; Si2024AST, ; Fan2023LargeLM, ). This opens an exciting direction for future research in Android performance analysis, promising to bridge the current gaps in existing techniques.
@@ 4.3. Constructing Benchmark Dataset is an Urgent Need
Similar to the shortage of tools, current datasets are too sporadic to reflect real performance issues, evidenced by the fact that only 24 (33.33%) of the 72 factors are covered by existing datasets. Even among the 24 factors, the data distribution is highly imbalanced, as the smallest dataset contains only five apps, while the largest has 324. In addition, the number of available data in each dataset is very small (usually only a few dozen examples), which is insufficient to support a fair evaluation of automated tools. Another critical issue is reproducibility. Our study indicates that only a small number of primary studies share valid datasets, making it difficult for future researchers to validate and evaluate tools. Furthermore, these datasets lack real-time updates for the latest app, thus containing outdated samples. Thus, it is urgent to construct up-to-date, comprehensive benchmark datasets to support fair comparisons, covering as many real-world cases as possible.
@@ 5. THREATS TO VALIDITY
Our study has several factors that could affect the validity of our findings in terms of large-scale exploration and literature review. For real-world data collection, the data filtering process may miss some relevant information. This is because the sentiment classification model (hartmann2023, ) used to filter user reviews may not guarantee 100% accuracy in identifying negative reviews. Additionally, the Google Play scraper might fail to capture all user reviews. Lastly, there could be biases in our manual checking process. To minimize subjective differences, two authors independently checked the filtered data. In cases of disagreement, they discussed and resolved whether to include the data.
For literature review, despite using a well-established methodology (Kitchenham2004ProceduresFP, ), we cannot guarantee that our empirical results cover all relevant works. To locate relevant studies, we formulated search strings from both Android and performance issues, as described in Section 3.1.2, and used five well-known electronic databases. To minimize the risk of missing relevant work, we conducted backward and forward snowballing and cross-checked our results.
@@ 6. Conclusion
In this paper, we conducted a comparative study to examine real-world Android performance issues encountered by users and developers, and to evaluate the extent to which current academic research addresses these practical challenges. By analyzing large-scale discussions from three popular crowd-sourced platforms on Android performance issues, alongside a review of 66 high-quality research papers, we identified several gaps between real-world challenges and the focus of academic research. Our main findings are as follows: (1) A taxonomy categorizes seven types of performance issues and 72 contributing factors. (2) Six common code patterns of these performance issues. (3) There is a misalignment in the primary concerns of researchers, developers, and users. (4) Only 42.86% of the real-world factors have been explored in academic research. (5) The majority of existing tools primarily utilize static analysis, dynamic analysis, and deep learning, with most tools detecting issues rather than fixing them. (6) Existing available tools target only 23.61% of identified factors. (7) Existing available datasets cover only 33.33% factors and lack sufficient data to support effective tool evaluation. Finally, we proposed future research directions to bridge the identified gaps.

### v3

@@ 3.2.2. Result - Technique
Lyu et al. (Lyu2018RemoveRF, ) 2018 206 Real-World
Energy Consumption
Responsiveness
Database Operations
Loop
 (Lyu2018RemoveRFTool, )
Vásquez et al. (LinaresVsquez2018MultiObjectiveOO, ) 2018 5 Real-World
Energy Consumption
Display Screen
 (LinaresVsquez2018MultiObjectiveOOTool, )
Habchi et al. (Habchi2019TheRO, ) 2019 324 Real-World
Memory Consumption
Energy Consumption
Responsiveness
UI Rendering
View
Inefficient Method
 (Habchi2019TheROData, )
Li et al. (Li2019CharacterizingAD, ) 2019 19 Real-World
Memory Consumption
Energy Consumption
Responsiveness
Image  (Li2019CharacterizingADTool, )
AsyncBench (Pan2020StaticAC, ) 2020 69 Micro
Memory Consumption
Energy Consumption
Responsiveness
AsyncTask  (Pan2020StaticACTool, )
Liu et al. (Liu2019DroidLeaksAC, ) 2020 34 Real-World
Memory Consumption
Energy Consumption
Resource
 (Liu2019DroidLeaksACTool, )
Das et al. (Das2020CharacterizingTE, ) 2020 316 Real-World
Memory Consumption
Energy Consumption
UI Rendering
Data Type
Inefficient Method
View
Wakelock
 (Das2020CharacterizingTETool, )
Lu et al. (Lu2022DetectingRU, ) 2022 35 Real-World
Memory Consumption
Energy Consumption
Resource
 (Lu2022DetectingRUTool, )
Janssen et al. (Janssen2022OnTI, ) 2022 40 Real-World
Energy Consumption
Responsiveness
UI Rendering
 (Janssen2022OnTITool, )
Petalotis et al. (Petalotis2024AnES, ) 2024 9 Real-World
Energy Consumption
Responsiveness
Advertisements
 (Petalotis2024AnESTool, )
Cruz et al. (Cruz2019CatalogOE, ) 2019 1,027 Real-World Energy Consumption
Display Screen
UI Rendering
Wakelock
Resource
Obsolete Task
 (Cruz2019CatalogOEDataset, )
Prestat et al. (Prestat2024DynAMICSAT, ) 2024 5 Real-World
Energy Consumption
Responsiveness
Memory Consumption
Wakelock
Inefficient Method
Resource
AsyncTask
View
 (Prestat2024DynAMICSATDataset, )
@@ 3.2.3. Result - Dataset
To collect datasets, several types of papers need to be included: 1) Tool papers that construct datasets to evaluate their designed tools; 2) Empirical studies that collect specific datasets to analyze performance issues, where labeled data can be used to support evaluation. After examining all the public links or websites within the 85 identified studies, we identified 12 (14.11%) unique, available, and open-sourced datasets, as shown in Table 4. Notably, we conducted the examination in Jun 2025. Since most of these datasets are designed for diverse performance factors, it becomes infeasible to ensure a fair comparison. In addition, the number of factors covered by these datasets is quite limited. Out of the 82 factors identified in our taxonomy, only 24 (29.27%) are addressed by existing datasets. This indicates a significant need for constructing more comprehensive datasets. Moreover, there is a considerable imbalance in dataset size. The smallest dataset contains only five apps, while the largest has 1,027, with an average of 174.1 apps per dataset. This imbalance makes it unrealistic for any single dataset to cover all potential issues comprehensively. Even if some tools are effective at detecting performance issues, they may still face soundness challenges due to false negatives. Therefore, more effort is needed to create datasets that are not only comprehensive but also more representative of the diverse range of performance issues.
Answers to RQ2.3: We identified 12 open-source datasets for evaluating performance issues; however, they encompass only 24 out of 82 factors (29.27%), highlighting a significant gap in tool evaluation. These datasets suffer from inadequate sample sizes, with the smallest containing only five apps and the largest consisting of 1,027.
@@ 4.1. Targeted Objective Imbalance
In Section 3.2.1, we found that there is a misalignment in the primary concerns of researchers, developers, and users: researchers primarily focus on energy consumption, developers prioritize memory consumption, while users are most concerned with responsiveness. This suggests that future research should increase focus on responsiveness and memory issues, as these have a more immediate impact on both end-users and developers. We also found two categories of performance issues, storage consumption and internet data usage, that remain unexplored in academic research, indicating potential areas for further investigation. Additionally, only 42.86% of the factors contributing to performance issues observed in real-world settings have been explored in academic research, remaining nearly 60% still awaiting for researchers to investigate. Interestingly, there are 19 factors among the 82 factors that lead to performance issues that have only been explored in literature but not observed in the real-world scenarios. This suggests that the issues currently discussed in academic research may not be of real concern to users or developers in real-world settings. Overall, this imbalance between academic research objectives and real-world challenges underscores the need for greater focus on the real-world performance issues and contributing factors that are most relevant to end-users and developers.
@@ 4.2. Automatic Performance Analysis Tools are Still in an Early Stage
Our study revealed that only 30 out of 82 factors (36.59%) are addressed by state-of-the-art (SOTA) tools, with most tools focusing on detection rather than resolution. In fact, only four out of the existing tools are dedicated to fixing performance issues. Although static, dynamic, and deep learning (DL) techniques have shown effectiveness in research experiments, applying these methods to holistically detect and fix performance issues in practice remains an unsolved challenge. Future research should focus not only on building comprehensive frameworks to detect real-world performance issues, but also on providing solutions for fixing these issues.
The lack of variety in technique selection hinders the soundness, as static approaches are not proficient in handling complicated code features (e.g., reflection (sun2021taming, ), obfuscation (samhi2022jucify, ), and hardening (Schloegel2022LokiHC, )), and it’s non-trivial for dynamic tools to achieve satisfactory results to trigger sufficient paths in practice. Thus, developing effective tools against performance issues will be a noteworthy research topic in the future. In fact, DL has made considerable achievements in software engineering (Wang2023MachineDeepLF, ; Feng2020CodeBERTAP, ; Gu2016DeepAL, ; Hu2018DeepCC, ; pan2024don, ), and these advanced techniques have demonstrated powerful capabilities in solving many complex tasks (i.e., code analysis (White2016DeepLC, ; Jebnoun2021ClonesID, ; Zhao2018DeepSimDL, ) and code generation (Wang2022TestDrivenML, ; zhang2025deployability, )). Applying advanced DL techniques to detect and fix Android performance issues is promising, although significant challenges remain in collecting a sufficient performance issue dataset for model training. The emergence of large language models (LLMs) also presents new possibilities. LLMs, with their in-context learning capabilities, could address the limitations of traditional DL approaches by enabling models to detect and resolve performance issues with fewer examples and greater adaptability (Hou2023LargeLM, ; Liao2023A3CodGenAR, ; Si2024AST, ; Fan2023LargeLM, ; pan2024large, ; han2024chase, ; Huang2022SEFK, ; Su2024EnhancingET, ). This opens an exciting direction for future research in Android performance analysis, promising to bridge the current gaps in existing techniques.
@@ 4.3. Constructing Benchmark Dataset is an Urgent Need
Similar to the shortage of tools, current datasets are too sporadic to reflect real performance issues, evidenced by the fact that only 24 (29.27%) of the 82 factors are covered by existing datasets. Even among the 24 factors, the data distribution is highly imbalanced, as the smallest dataset contains only five apps, while the largest has 1,027. In addition, the amount of available data in each dataset is very small (usually only a few dozen examples), which is insufficient to support a fair evaluation of automated tools. Another critical issue is reproducibility. Our study indicates that only a small number of primary studies share valid datasets, making it difficult for future researchers to validate and evaluate tools. Furthermore, these datasets lack real-time updates for the latest app, thus containing outdated samples. Thus, it is urgent to construct up-to-date, comprehensive benchmark datasets to support fair comparisons, covering as many real-world cases as possible.
@@ 5. Threats to Validity
Our study is subject to several potential threats to validity, which we categorize into internal and external aspects.
5.1. Internal Validity
@@
This concerns the soundness of our data collection and literature review processes. For real-world data filtering, the process may miss some relevant information. This is because the sentiment classification model (hartmann2023, ) used to filter user reviews may not guarantee 100% accuracy in identifying negative reviews. Additionally, the Google Play scraper might fail to capture all user reviews. Lastly, there could be biases in our manual checking process. To minimize subjective differences, two authors independently checked the filtered data. In cases of disagreement, they discussed and resolved whether to include the data. For literature review, despite using a well-established methodology (Kitchenham2004ProceduresFP, ), we cannot guarantee that our empirical results cover all relevant works. To locate relevant studies, we formulated search strings from both Android and performance issues, as described in Section 3.1.2, and used five well-known electronic databases. To minimize the risk of missing relevant work, we conducted backward and forward snowballing and cross-checked our results.
5.2. External Validity
This concerns the generalizability of our findings beyond the studied datasets and contexts. Regarding the representativeness of data sources, our real-world data was derived from open-source and community-driven platforms such as GitHub, Stack Overflow, and Google Play. While these platforms reflect commonly reported issues from general developers and users, they may not capture the performance concerns of proprietary or enterprise-level applications developed by organizations such as Google or Meta. Furthermore, Stack Overflow discussions may be biased toward novice or intermediate developers, as experienced professionals may participate less frequently. As for the research scope, this study focuses exclusively on performance issues in Android mobile applications. Other Android-based platforms, such as Android TV, Android Auto, or IoT systems, may exhibit different performance behaviors due to variations in hardware, user interactions, and platform-specific constraints. These platforms lie beyond the scope of this work and represent promising avenues for future research. Another external threat lies in the lack of user-centric validation. While our developer survey provides evidence that the taxonomy reflects practical development concerns, we did not conduct a complementary study with end users. Consequently, our taxonomy may not fully capture performance issues that users subjectively perceive as most critical. Future work could incorporate user-centric validation, for example through controlled user studies or large-scale questionnaires, to further strengthen the external validity of our findings.
6. Related Work
To investigate Android performance issues, numerous studies have been conducted, which can be broadly categorized into two directions: (1) empirical studies based on real-world data sources, and (2) literature reviews that summarize the current state of performance-related research.
6.1. Empirical Studies on Android Performance Issues
Numerous studies have explored Android performance issues using real-world sources (Kumari2024AnES, ; Noor2022EndUP, ; Das2016AQA, ; Vsquez2015HowDD, ; Liu2014CharacterizingAD, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Habchi2019TheRO, ; Habchi2020AndroidCS, ; MazueraRozo2020InvestigatingTA, ; Prestat2022AnES, ). These studies typically focus on a single type of data source and adopt either the user or developer perspective. For example, Noor et al. (Noor2022EndUP, ) focused on the user perspective by examining 368,704 low-rated Google Play reviews from 55 apps. They identified eight categories of performance issues reported by users, including responsiveness delays, graphical glitches, and image rendering problems. Kumari et al. (Kumari2024AnES, ) analyzed 385 Stack Overflow posts to understand performance issues discussed by developers. They proposed two taxonomies: one categorizing eight performance issue types (e.g., UI/UX, graphics, and development practices), and another describing their associated causes. Similarly, Das et al. (Das2016AQA, ), Vásquez et al. (Vsquez2015HowDD, ) and Liu et al. (Liu2014CharacterizingAD, ) analyzed GitHub issue reports and commit histories to categorize performance problems that developers frequently encounter.
While these studies effectively identify performance issues and root causes from user and developer perspectives, they do not assess whether these real-world concerns are adequately addressed by academic research. Moreover, although some works propose taxonomies, the classification dimensions and terminology vary significantly, hindering comparison and integration across studies. In contrast, our study integrates data from multiple sources, including Google Play (user perspective), Stack Overflow and GitHub (developer perspective), and academic literature (researcher perspective). Building on this integration, we adopt a hierarchical consequence–cause structure that explicitly links observable performance issues to their underlying root causes, enabling direct mapping to existing tools and datasets for comprehensive gap analysis.
6.2. Literature Reviews on Android Performance Issues
To assess the current state of Android performance analysis, several systematic literature reviews have summarized the performance-related issues studied in academic research (Wu2023ASL, ; Fawad2024AndroidSC, ; Hort2021ASO, ). For example, Wu et al. (Wu2023ASL, ) reviewed 35 studies on Android code smells and identified five major categories, including performance-related smells such as energy inefficiency and memory usage. Similarly, Fawad et al. (Fawad2024AndroidSC, ) analyzed 79 papers and extracted 237 types of Android code smells, covering concerns such as durable wakelocks, member-ignoring methods, and inefficient database access, many of which are related to performance. In addition, Hort et al. (Hort2021ASO, ) reviewed 156 papers on non-functional performance optimization in Android apps. These studies focused on four key performance characteristics (e.g., responsiveness, launch time, memory consumption, and energy usage), and summarized common optimization strategies such as anti-pattern removal, refactoring, and preloading.
While these reviews offer valuable insights into performance issues and proposed solutions, they do not evaluate whether academic efforts effectively address the most critical challenges in real-world development. They also overlook the extent to which existing tools and datasets support the detection and evaluation of such issues. Consequently, the actual alignment between academic research and practical concerns remains unclear. In contrast, our study first identifies the full spectrum of real-world performance issues. We then analyze how many of these issues are examined in the literature, which tools are available to address them, and whether corresponding datasets exist to support empirical evaluation. This enables a comprehensive understanding of research coverage and highlights gaps to guide future work.
@@ 7. Conclusion
In this paper, we conducted a comparative study to examine real-world Android performance issues encountered by users and developers, and to evaluate the extent to which current academic research addresses these practical challenges. By analyzing large-scale discussions from three popular crowd-sourced platforms on Android performance issues, alongside a review of 85 high-quality research papers, we identified several gaps between real-world challenges and the focus of academic research. Our main findings are as follows: (1) A taxonomy categorizes seven types of performance issues and 82 contributing factors. (2) Six common code patterns of these performance issues. (3) There is a misalignment in the primary concerns of researchers, developers, and users. (4) Only 42.86% of the real-world factors have been explored in academic research. (5) The majority of existing tools primarily utilize static analysis, dynamic analysis, and deep learning, with most tools detecting issues rather than fixing them. (6) Existing available tools target only 36.59% of identified factors. (7) Existing available datasets cover only 29.27% of factors and lack sufficient data to support effective tool evaluation. Finally, we proposed future research directions to bridge the identified gaps.